# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ Abstract.
The increasing adoption of Large Language Models (LLMs) in software engineering has sparked interest in their use for software vulnerability detection. However, the rapid development of this field has resulted in a fragmented research landscape, with diverse studies that are difficult to compare due to differences in, e.g., system designs and dataset usage. This fragmentation makes it difficult to obtain a clear overview of the state-of-the-art or compare and categorize studies meaningfully. In this work, we present a comprehensive systematic literature review (SLR) of LLM-based software vulnerability detection. We analyze 227 studies published between January 2020 and June 2025, categorizing them by task formulation, input representation, system architecture, and adaptation techniques. Further, we analyze the datasets used, including their characteristics, vulnerability coverage, and diversity. We present a fine-grained taxonomy of vulnerability detection approaches, identify key limitations, and outline actionable future research opportunities. By providing a structured overview of the field, this review improves transparency and serves as a practical guide for researchers and practitioners aiming to conduct more comparable and reproducible research. We publicly release all artifacts and maintain a living repository of LLM-based software vulnerability detection studies.
@@ 1. Introduction
Vulnerability assessment pipelines as found in industrial development workflows typically consist of multiple steps: detection (identifying whether code contains a vulnerability), localization (pinpointing the exact location in the code), severity estimation (evaluating potential impact), and repair (modifying the code to mitigate the vulnerability), often initiated reactively when a new CVE is disclosed. Despite advances in tooling, e.g., static analysis tools, these processes remain largely manual or only partially automated due to the reliance on expert-defined rule sets and high false positive rates of such tools (Croft et al., 2023). As such, existing workflows are insufficient to meet the scale of remediation demands.
@@
At the same time, the adoption of Large Language Models (LLMs)111Due to the lack of a formal definition for LLMs, in this work, we use the term LLM broadly to encompass both large-scale language models (e.g., GPT-4) and smaller pre-trained models such as BERT and CodeBERT (Zhao et al., 2024b). for software engineering tasks has grown rapidly since the introduction of BERT (Devlin et al., 2019) in 2018. LLMs are deep neural network models, typically based on transformer architectures (Vaswani et al., 2017), that are pre-trained on massive corpora of natural language. LLMs have demonstrated impressive capabilities in understanding and generating code. Industry forecasts further emphasize the transformative potential of LLMs. For instance, Anthropic CEO Dario Amodei predicts that in mid to end 2025, AI is already writing 90% of the code, with developers primarily specifying design goals and constraints (Amodei, 2025).
@@
However, while LLMs hold promise in code generation, they also introduce new risks. Generated code often lacks awareness of existing libraries or internal codebases, which may lead to redundancies and the creation of syntactically different but semantically equivalent code. This complicates the detection of vulnerabilities, especially when they are subtle or semantically nuanced. Additionally, LLMs are prone to hallucinations and the generation of insecure code (Pearce et al., 2025; Liu et al., 2024d, b; Fu et al., 2025). Development teams may be unaware of vulnerabilities introduced by generated code, whether in their codebase, third-party components, or libraries.
@@
The combination of LLM-generated code and the increasing volume of vulnerabilities calls for automated, scalable, and reliable vulnerability assessment pipelines. Recent research has begun to explore the use of LLMs for vulnerability remediation. In a prior study (Kaniewski et al., 2024), we surveyed early work on LLM-based vulnerability handling, including detection, localization, and repair, and highlighted open challenges with regard to scalability, data diversity, and model reliability. We continued to extensively document the progress in this field, with a focus on vulnerability detection studies as a prerequisite for all subsequent remediation steps. As the number of studies on LLM-based vulnerability detection continues to grow, we observe a wide variety of system architectures, adaptation techniques, and evaluation methodologies. This heterogeneity leads to two pressing challenges: First, there is currently no comprehensive survey that maps the landscape of LLM-based software vulnerability detection methods, their system designs, and dataset usage. As a result, researchers face difficulties identifying trends, gaps, or best practices in this rapidly evolving area. Second, studies frequently adopt custom datasets, metrics, and data splits, making it difficult to assess progress, reproduce results, or compare models fairly. To address these challenges, we contribute this systematic literature review (SLR) of LLM-based software vulnerability detection studies. Specifically, the contributions are as follows:
@@
We conduct a SLR of studies covering LLM-based software vulnerability detection in the time from January 2020 to June 2025, covering 227 studies.
@@
We present a comprehensive taxonomy that categorizes existing approaches according to task formulation, input representation, system architecture, adaptation technique, and dataset usage. This taxonomy allows for meaningful comparison of studies.
@@
The remainder of this paper is structured as follows: We review related surveys in Section 2, and describe the SLR methodology in Section 3. Section 4 presents the comprehensive taxonomy for LLM-based vulnerability detection. We extend this taxonomy and analyze datasets used in Section 5. We discuss current limitations and future research opportunities in Section 6. Finally, we conclude this review in Section 7.
2. Related Surveys
@@
With the growing adoption of LLMs across various domains, numerous works have emerged to explore their applications. We review related surveys that intersect most closely with the topic of LLM-based software vulnerability detection.
| Coverage | |||||||
| Domain and Reference | Published | Time Frame | #Studies | LLMs | Input Repr. | Adapt. Techniques | Datasets |
| Cybersecurity | |||||||
| Xu et al. (Xu et al., 2025) | Jul 2024 | 2020 – 2024 | 127 (17) |
| Zhang et al. (Zhang et al., 2025a) | Dec 2024 | 2023 – 2024 | 300 (29) |
| Ferrag et al. (Ferrag et al., 2025a) | Jan 2025 | n/s | n/s (6) |
| Software Engineering | |||||||
| Zhang et al. (Zhang et al., 2024b) | Sep 2024 | 2020 – 2024 | 947 (71) |
| Hou et al. (Hou et al., 2024) | Dec 2024 | 2017 – 2024 | 395 (18) |
| Hu et al. (Hu et al., 2025) | May 2025 | 2007 – 2024 | 191 (30) |
| Vulnerability Detection | |||||||
| Shiri Harzevili et al. (Shiri Harzevili et al., 2024) | Nov 2024 | 2011 – 2024 | 138 (4) |
| Shereen et al. (Shereen et al., 2024) | Dec 2024 | 2006 – 2024 | 79 (10) |
| Shimmi et al. (Shimmi et al., 2025) | Jun 2025 | 2018 – 2023 | 98 (8) |
| Taghavi Far and Feyzi (Taghavi Far and Feyzi, 2025) | Jan 2025 | n/s | 119 (n/s) |
| Basic and Giaretta (Basic and Giaretta, 2025) | Apr 2025 | n/s | n/s (n/s) |
| Zhou et al. (Zhou et al., 2025a) | Oct 2024 | 2018 – 2024 | 58 (40) |
| Sheng et al. (Sheng et al., 2025) | Feb 2025 | 2019 – 2024 | 58 |
| This survey | July 2025 | 2020 – 2025 | 227 |
In Table 1, we provide an overview of related surveys, with particular attention to their coverage of input representations, adaptation techniques, and datasets. Several surveys analyze the application of LLMs in the domain of cybersecurity (Xu et al., 2025; Zhang et al., 2025a; Ferrag et al., 2025a), typically addressing vulnerability detection as one of several sub-tasks. Similarly, surveys in the domain of software engineering (Hou et al., 2024; Zhang et al., 2024b; Hu et al., 2025) categorize LLM usage by phases of the software development life cycle. Vulnerability detection is typically a part of software testing (Zhang et al., 2024b) or software quality assurance (Hou et al., 2024). Consequently, these broader surveys provide limited insights into LLM applications specific to software vulnerability detection, such as task-specific adaptation techniques and dataset usage. Shiri Harzevili et al. (Shiri Harzevili et al., 2024), Shereen et al. (Shereen et al., 2024), and Shimmi et al. (Shimmi et al., 2025) focus on the application of machine learning and deep learning techniques for software vulnerability detection. Their discussion of studies applying LLMs is limited, covering only four, ten, and eight studies, respectively. Taghavi Far and Feyz (Taghavi Far and Feyzi, 2025) do not focus specifically on software vulnerability detection but consider a wider scope, including phishing detection, threat detection in logs, and patch recommendation. The surveys by Basic and Giaretta (Basic and Giaretta, 2025) and Zhou et al. (Zhou et al., 2025a) are more closely aligned with LLM-based software vulnerability detection. However, their scope extends to code security and vulnerability repair, providing only limited insights specific to software vulnerability detection. Sheng et al. (Sheng et al., 2025) present an overview of LLMs applied to vulnerability detection, datasets, metrics, and techniques, focusing on code processing and prompt engineering methods. In contrast to prior surveys, this SLR provides a focused and up-to-date overview of 227 studies, presenting a comprehensive taxonomy of LLM-based software vulnerability detection. We classify existing work across detection tasks and objectives, input representations, system architectures, and adaptation techniques. This is the first comprehensive taxonomy for the research area of LLM-based software vulnerability detection, systematically analyzing state-of-the-art techniques.
@@
A further distinguishing aspect of this SLR is its in-depth analysis of vulnerability detection datasets. High-quality, realistic, and diverse datasets play a crucial role in learning-based software vulnerability detection, prompting studies to analyze, classify, and evaluate their characteristics, quality, and impact on model performance. Previous surveys have touched upon datasets: Shiri Harzevili et al. (Shiri Harzevili et al., 2024) categorize dataset sources into benchmark, hybrid, open source software, and repository. Shereen et al. (Shereen et al., 2024) compare frequently used datasets with respect to properties such as language, size, annotation, and granularity. Similarly, Shimmi et al. (Shimmi et al., 2025) list used datasets with respect to granularity and language. Ferrag et al. (Ferrag et al., 2025a) provide an overview of datasets suitable for fine-tuning LLMs in security applications. Hou et al. (Hu et al., 2025) analyze benchmarks for evaluating LLM-based solutions, focusing on construction methods, programming languages, and metrics. Sheng et al. (Sheng et al., 2025) discuss datasets used for LLM-based vulnerability detection, focusing on granularity and programming languages. In contrast to previous surveys that primarily list or categorize datasets by properties such as granularity, language, or source, this review offers a more comprehensive and structured analysis. We categorize datasets using a detailed taxonomy covering type, granularity, source, and labeling methodology, and investigate additional quality dimensions, such as class balance, CWE diversity, and distribution, offering an in-depth analysis of represented vulnerabilities. Further, we discuss emerging trends and use cases in dataset utilization.
@@
Other related studies focus explicitly on vulnerability detection datasets: Lin et al. (Lin et al., 2022b) examine dataset construction methodologies, comparing datasets across six dimensions, such as granularity, vulnerability type, and labeling approach. Guo et al. (Guo et al., 2024a) analyze selected datasets in terms of vulnerability distributions and types. Jain et al. (Jain et al., 2023) provide a code-centric evaluation of commonly used C/C++datasets for deep learning-based vulnerability detection. Moreover, the works (Guo and Bettaieb, 2023; Croft et al., 2023; Risse et al., 2025; Yadav and Wilson, 2024) provide in-depth analyses of quality issues in few selected vulnerability datasets, such as data imbalance, low vulnerability coverage, biased vulnerability distribution, high duplication rates, and label noise. Despite this growing body of work, a targeted and comprehensive overview of vulnerability detection datasets commonly used for LLM-based software vulnerability detection approaches, along with their limitations and suitability remains missing. This survey addresses this gap, offering a taxonomy and systematic analysis of datasets, their CWE coverage, diversity, and usage, thereby facilitating better comparability and benchmarking in future research.

### v2

@@ Abstract.
The increasing adoption of Large Language Models (LLMs) in software engineering has sparked interest in their use for software vulnerability detection. However, the rapid development of this field has resulted in a fragmented research landscape, with diverse studies that are difficult to compare due to differences in, e.g., system designs and dataset usage. This fragmentation makes it difficult to obtain a clear overview of the state-of-the-art or compare and categorize studies meaningfully. In this work, we present a comprehensive systematic literature review (SLR) of LLM-based software vulnerability detection. We analyze 263 studies published between January 2020 and November 2025, categorizing them by task formulation, input representation, system architecture, and techniques. Further, we analyze the datasets used, including their characteristics, vulnerability coverage, and diversity. We present a fine-grained taxonomy of vulnerability detection approaches, identify key limitations, and outline actionable future research opportunities. By providing a structured overview of the field, this review improves transparency and serves as a practical guide for researchers and practitioners aiming to conduct more comparable and reproducible research. We publicly release all artifacts and maintain a living repository of LLM-based software vulnerability detection studies at https://github.com/hs-esslingen-it-security/Awesome-LLM4SVD.
@@ 1. Introduction
Vulnerability assessment pipelines as found in industrial development workflows typically consist of multiple steps: detection (identifying whether code contains a vulnerability), localization (pinpointing the exact location in the code), severity estimation (evaluating potential impact), and repair (patching the code to mitigate the vulnerability), often initiated reactively when a new CVE is disclosed. Despite advances in tooling, e.g., static application security testing (SAST) tools, these processes remain largely manual or only partially automated due to the dependency on expert-defined rules (cf. Taghavi and Feyzi (Taghavi Far and Feyzi, 2025) for a comparison between LLMs and established software vulnerability detection methods). Further, these tools suffer from high false positive rates (Croft et al., 2023). As such, existing workflows are insufficient to meet the scale of remediation demands.
@@
At the same time, the adoption of Large Language Models (LLMs)111Due to the lack of a formal definition for LLMs, in this work, we use the term LLM broadly to encompass both large-scale language models (e.g., GPT-4) and smaller pre-trained models, such as BERT and CodeBERT (Zhao et al., 2024b). We differentiate architectures and scales of LLMs later on in the body of this work. for software engineering tasks has grown rapidly since the introduction of BERT (Devlin et al., 2019) in 2018. LLMs are deep neural network models, typically based on Transformer architectures (Vaswani et al., 2017), that are pre-trained on massive corpora of natural language. LLMs have demonstrated impressive capabilities in understanding and generating code. Industry forecasts further emphasize the transformative potential of LLMs in code generation (Amodei, 2025).
@@
However, while LLMs hold promise in code generation, they also introduce new risks. Generated code often lacks awareness of existing libraries or internal codebases, which may lead to redundancies and the creation of syntactically different but semantically equivalent code. This complicates the detection of vulnerabilities, especially when they are subtle or semantically nuanced. Additionally, LLMs are prone to hallucinations and the generation of insecure code (Pearce et al., 2025; Liu et al., 2024d, c; Fu et al., 2025a). Development teams may be unaware of vulnerabilities introduced by generated code, whether in their codebase, third-party components, or libraries.
@@
The combination of LLM-generated code and the increasing volume of vulnerabilities calls for automated, scalable, and reliable vulnerability assessment pipelines. LLMs’ ability to generalize across programming languages and representations, as well as their high flexibility through prompt adaptation, make them a promising approach for software vulnerability assessment. Therefore, recent research has begun to explore the use of LLMs. In a prior study (Kaniewski et al., 2024), we surveyed early work on LLM-based vulnerability handling, including detection, localization, and repair, and highlighted open challenges with regard to scalability, data diversity, and model reliability. We continued to extensively document the progress in this field, with a focus on vulnerability detection studies as a prerequisite for all subsequent remediation steps. As the number of studies on LLM-based vulnerability detection continues to grow, we observe a wide variety of system architectures, adaptation techniques, and evaluation methodologies. This heterogeneity leads to two pressing challenges: First, there is currently no comprehensive survey that maps the landscape of LLM-based software vulnerability detection methods, their system designs, and dataset usage. As a result, researchers face difficulties identifying trends, gaps, or best practices in this rapidly evolving area. Second, studies frequently adopt custom datasets, metrics, and data splits, making it difficult to assess progress, reproduce results, or compare models fairly. To address these challenges, we contribute this systematic literature review (SLR) of LLM-based software vulnerability detection studies. Specifically, our contributions are as follows:
@@
We conduct a SLR of studies covering LLM-based software vulnerability detection in the time from January 2020 to November 2025, covering 263 studies.
@@
We present a comprehensive taxonomy that categorizes existing approaches based on task formulation (classification and generation), input representation (code and auxiliary information), system architecture (LLM usage), adaptation and orchestration technique, as well as dataset usage. This taxonomy allows for meaningful comparison of studies.
@@
The remainder of this paper is structured as follows: We review related surveys in Section 2, and describe the SLR methodology in Section 3. Section 4 presents the comprehensive taxonomy for LLM-based vulnerability detection. We extend this taxonomy and analyze datasets used in Section 5. We discuss current limitations and future research opportunities in Section 6. Finally, we address limitations of this SLR in Section 7 and conclude this review in Section 8.
2. Related Work
@@
With the growing adoption of LLMs across various domains, numerous works have emerged to explore their applications. In this section, we discuss related works, covering related SLRs and surveys that intersect most closely with the topic of LLM-based software vulnerability detection, as well as works focusing on vulnerability datasets.
2.1. Related SLRs and Surveys
In Table 1, we provide an overview of related surveys, with particular attention to their coverage of the main aspects of this SLR, i.e., focus on LLM-based studies, input representations, adaptation techniques, analysis of datasets, and a comprehensive taxonomy. In the following, we discuss these surveys and how this SLR differs in coverage from previous works.
Several surveys analyze the application of LLMs in the domain of cyber security (Xu et al., 2025; Ferrag et al., 2025a), addressing vulnerability detection as one of several sub-tasks. Xu et al. (Xu et al., 2025), for example, cover the domains of software and system security, network security, information and content security, hardware security, and blockchain security. The authors focus on what LLMs have been employed to support these security tasks, and discuss domain-specific adaptation techniques, i.e., fine-tuning, prompt engineering, and external augmentation. Further, they address dataset collection and pre-processing techniques. Ferrag et al. (Ferrag et al., 2025a) focus on LLMs employed for cyber security and their optimization strategies, briefly discussing advanced training and fine-tuning techniques. The authors also provide an overview of software vulnerability datasets used for fine-tuning.
Surveys in the domain of software engineering (Hou et al., 2024; Zhang et al., 2024a; Fu et al., 2025b) categorize LLM usage by phases of the software development life cycle. Vulnerability detection is typically part of software testing (Zhang et al., 2024a) or software quality assurance (Hou et al., 2024). Fu et al. (Fu et al., 2025b) focus on AI for DevSecOps, covering software vulnerability detection and classification. Due to the broader scope of AI, the survey covers fewer studies applying LLM. Zhang et al. (Zhang et al., 2024a) focus on the LLM perspective, providing a comprehensive overview of code LLMs, discussing pre-training tasks, and downstream applications. Hou et al. (Hou et al., 2024) discuss techniques used to optimize LLMs, including prompt engineering and fine-tuning techniques. The authors also provide an overview of input forms used for diverse software engineering tasks (i.e., token- and tree/graph-based) and highlight data collection, selection, and pre-processing processes. While these surveys provide overviews of the application of LLMs in broader domains, we focus on insights into the state-of-the-art adaptation and orchestration techniques, input representations, and datasets that are specific to software vulnerability detection.
| Coverage | ||||||||||
| Reference | Published | Focus | Time Frame | #Studies (all — SVD) |
SLR |
LLM Focus |
Input Repres. |
Adaptation Techniques |
Dataset Analysis |
Taxonomy |
| Cyber Security | ||||||||||
| Xu et al. (Xu et al., 2025) | Sep 2025 | Cyber security | 2020-2025 | 185 — 22 | ✓ | ● | ○ | ● | ◐ | ○ |
| Ferrag et al. (Ferrag et al., 2025a) | May 2025 | Cyber security | n/s | n/s — 9 | ✗ | ● | ○ | ◐ | ◐ | ○ |
| Software Engineering | ||||||||||
| Zhang et al. (Zhang et al., 2024a) | Sep 2024 | Software engineering | 2020-2024 | 947 — 71 | ✓ | ● | ○ | ○ | ○ | ○ |
| Hou et al. (Hou et al., 2024) | Dec 2024 | Software engineering | 2017-2024 | 395 — 18 | ✓ | ● | ● | ● | ◐ | ○ |
| Fu et al. (Fu et al., 2025b) | Apr 2025 | DevSecOps | 2017-2023 | 99 — 9 | ✓ | ○ | ○ | ○ | ◐* | ○ |
| Vulnerability Detection | ||||||||||
| Shiri Harzevili et al. (Shiri Harzevili et al., 2024) | Nov 2024 | Autom. software vulnerability detection | 2011-2024 | 138 — 7 | ✓ | ○ | ● | ○ | ◐ | ○ |
| Shereen et al. (Shereen et al., 2024) | Dec 2024 | Autom. vulnerability detection | 2006-2024 | 79 — 10 | ✗ | ○ | ◐ | ○ | ◐* | ○ |
| Shimmi et al. (Shimmi et al., 2025) | Jun 2025 | Software vulnerability detection | 2018-2023 | 98 — 13 | ✓ | ○ | ● | ○ | ◐* | ● |
| Taghavi Far and Feyzi (Taghavi Far and Feyzi, 2025) | Feb 2025 | Software vulnerability detection | n/s | n/s — n/s | ✓ | ● | ○ | ◐ | ◐ | ○ |
| Basic and Giaretta (Basic and Giaretta, 2025) | Apr 2025 | Code security | n/s | n/s — n/s | ✓ | ● | ○ | ◐ | ○ | ○ |
| Zhou et al. (Zhou et al., 2025a) | May 2025 | Vulnerability detection and repair | 2018-2024 | 58 — 40 | ✓ | ● | ○ | ● | ○ | ○ |
| Germano et al. (Germano et al., 2025) | Nov 2025 | Code vulnerability detection, repair, expl. | 2018-2025 | 208 — 190 | ✓ | ● | ○ | ○ | ◐* | ○ |
| Sheng et al. (Sheng et al., 2025) | Sep 2025 | Software security, vulnerability detection | 2017-2025 | 58 — 58 | ✗ | ● | ◐ | ● | ◐ | ○ |
| *List only | ||||||||||
| This SLR | Dec 2025 | Software vulnerability detection | 2020-2025 | 263 | ✓ | ● | ● | ● | ● | ● |
Shiri Harzevili et al. (Shiri Harzevili et al., 2024), Shereen et al. (Shereen et al., 2024), and Shimmi et al. (Shimmi et al., 2025) focus on the application of machine learning and deep learning techniques for vulnerability detection, discussing seven, ten, and thirteen studies applying LLMs, respectively. These studies discuss machine learning and deep learning specific input (tree, graph, token) (Shiri Harzevili et al., 2024; Shimmi et al., 2025) or program representations (source, binary, intermediate) (Shereen et al., 2024) and provide an overview of common datasets. Shimmi et al. (Shimmi et al., 2025) also provide a taxonomy of methods, feature representations, and embedding techniques. In contrast to these surveys, we focus explicitly on LLM-based studies, discussing and categorizing LLM-specific input representations and adaptation techniques, as well as challenges with existing datasets and dataset usage. Taghavi Far and Feyz (Taghavi Far and Feyzi, 2025) consider a broader scope of software vulnerability detection, including, e.g., phishing detection, threat detection, and patch recommendation. The authors compare the application of LLMs to traditional program analysis techniques, discuss various fine-tuning techniques, and present selected pre-training and fine-tuning datasets. The surveys by Basic and Giaretta (Basic and Giaretta, 2025), Zhou et al. (Zhou et al., 2025a), and Germano et al. (Germano et al., 2025) are closely aligned with software vulnerability detection, extending to code security and vulnerability repair. Specifically, Basic and Giaretta (Basic and Giaretta, 2025) focus on vulnerabilities introduced by LLM-generated code as well as LLMs for detecting and fixing vulnerabilities (in a unified process), discussing various prompting strategies. Zhou et al. (Zhou et al., 2025a) discuss LLMs for vulnerability detection and repair, focusing on early fine-tuning, prompt engineering (zero-shot, few-shot), and retrieval augmentation techniques. Germano et al. (Germano et al., 2025) survey studies for vulnerability detection, repair, and explanation with LLMs, discussing the different models and programming languages applied as well as the distribution of datasets. Most closely to this SLR is the survey by Sheng et al. (Sheng et al., 2025), covering vulnerability detection techniques and insights. The authors discuss code pre-processing techniques, cover prompt engineering as well as fine-tuning techniques, and provide an overview of languages and granularities in benchmarks and datasets. In contrast, this SLR further offers a fine-grained taxonomy of prompting and fine-tuning techniques, incorporating learning paradigms and orchestration strategies. Furthermore, we introduce a temporal analysis of dataset usage and critically evaluate the vulnerability diversity and coverage of common benchmarks. Finally, we deepen the discussion on input representations and the role of auxiliary information in both prompting and tuning contexts, as well as classification and generation tasks.
Overall, this SLR extends the field through a dedicated and in-depth overview of 263 studies on LLM-based software vulnerability detection. The further this research field evolves, the more patterns and divergences can be identified. By systematically classifying studies using a comprehensive taxonomy, this SLR reveals focused insights that remain hidden in broader surveys, including how vulnerabilities are represented, which (new) techniques have been adapted and are effective for detection, and how study design influences applicability and comparability.
2.2. Vulnerability Datasets
@@
A further distinguishing aspect of this SLR from previous surveys is its in-depth analysis of vulnerability detection datasets. High-quality, realistic, and diverse datasets play a crucial role in learning-based software vulnerability detection, prompting studies to analyze, classify, and evaluate their characteristics, quality, and impact on model performance. Discussed surveys have touched upon vulnerability detection datasets: Xu et al. (Xu et al., 2025) and Hou et al. (Hou et al., 2024) categorize dataset collection, i.e., open-source, collected, constructed, or industrial, as well as dataset type, i.e., code-based, text-based, graph-based, software repository-based, or combined. Shiri Harzevili et al. (Shiri Harzevili et al., 2024) categorize dataset sources into benchmark, hybrid, open source software, and repository. Further, they differentiate code, text, numerical, and hybrid dataset types. Shereen et al. (Shereen et al., 2024) compare frequently used datasets with respect to properties such as language, size, annotation, and granularity. Similarly, other discussed surveys (Ferrag et al., 2025a; Fu et al., 2025b; Shimmi et al., 2025; Taghavi Far and Feyzi, 2025; Germano et al., 2025; Sheng et al., 2025) list used datasets with respect to, e.g., size, programming language, granularity, or source. In contrast to previous surveys that primarily list datasets, we present a taxonomy and comprehensive analysis. We categorize datasets along type, granularity, source, and labeling methodology, and investigate additional quality dimensions, such as class balance, CWE diversity, and distribution, offering an in-depth analysis of represented vulnerabilities. Further, we add a time component and discuss emerging trends and use cases for datasets in LLM-based software vulnerability detection.
@@
Other related studies focus explicitly on vulnerability detection datasets: Hu et al. (Hu et al., 2025) analyze benchmarks for evaluating LLM-based solutions, focusing on construction methods, programming languages, and metrics. Lin et al. (Lin et al., 2022b) examine dataset construction methodologies, comparing datasets across six dimensions, such as granularity, vulnerability type, and labeling approach. Guo et al. (Guo et al., 2024a) analyze selected datasets in terms of vulnerability distributions and types. Jain et al. (Jain et al., 2023) provide a code-centric evaluation of commonly used C/C++datasets for deep learning-based vulnerability detection. Moreover, the works (Guo and Bettaieb, 2023; Croft et al., 2023; Risse et al., 2025; Yadav and Wilson, 2024) provide in-depth analyses of quality issues in selected vulnerability datasets, such as data imbalance, low vulnerability coverage, biased vulnerability distribution, high duplication rates, and label noise. Despite this growing body of work, a targeted and comprehensive overview of vulnerability detection datasets commonly used for LLM-based software vulnerability detection approaches, along with their limitations and suitability, remains missing. This survey addresses this gap, offering a taxonomy and systematic analysis of datasets, their CWE coverage, diversity, and usage, thereby facilitating better comparability and benchmarking in future research.